# Unifying Large Language Models and Knowledge Graphs: A Roadmap

### v1

@@ 4.3.1 KGs for LLM Probing
However, LAMA ignores the fact that the prompts are inappropriate. For example, the prompt “Obama worked as a _” may be more favorable to the prediction of the blank by the language models than “Obama is a _ by profession”. Thus, LPAQA [124] proposes a mining and paraphrasing-based method to automatically generate high-quality and diverse prompts for a more accurate assessment of the knowledge contained in the language model. Moreover, Adolphs et al. [126] attempt to use examples to make the language model understand the query, and experiments obtain substantial improvements for BERT-large on the T-REx data. Unlike using manually defined prompt templates, Autoprompt [125] proposes an automated method, which is based on the gradient-guided search to create prompts.
@@ 4.3.2 KGs for LLM Analysis
Tasks Method Year Technique
LLM-augmented KG embedding Pretrain-KGE [96] 2020 LLMs as Text Encoders
KEPLER [40] 2020 LLMs as Text Encoders
Nayyeri et al. [136] 2022 LLMs as Text Encoders
Huang et al. [137] 2022 LLMs as Text Encoders
CoDEx [138] 2022 LLMs as Text Encoders
LMKE [139] 2022 LLMs for Joint Text and KG Embedding
kNN-KGE [140] 2022 LLMs for Joint Text and KG Embedding
LambdaKG [141] 2023 LLMs for Joint Text and KG Embedding
LLM-augmented KG completion KG-BERT [26] 2019 Joint Encoding
MTL-KGC [142] 2020 Joint Encoding
PKGC [143] 2022 Joint Encoding
LASS [144] 2022 Joint Encoding
MEM-KGC [145] 2021 MLM Encoding
OpenWorld KGC [146] 2023 MLM Encoding
StAR [147] 2021 Separated Encoding
SimKGC [148] 2022 Separated Encoding
LP-BERT [149] 2022 Separated Encoding
GenKGC [98] 2022 LLM as decoders
KGT5 [150] 2022 LLM as decoders
KG-S2S [151] 2022 LLM as decoders
AutoKG [95] 2023 LLM as decoders
LLM-augmented KG construction LDET [152] 2019 Entity Typing
BOX4Types [153] 2021 Entity Typing
LRN [154] 2021 Entity Linking
TempEL [155] 2023 Entity Linking
BertCR [156] 2019 CR (Within-document)
Spanbert [157] 2020 CR (Within-document)
CDLM [158] 2021 CR (Cross-document)
CrossCR [159] 2021 CR (Cross-document)
CR-RL [160] 2021 CR (Cross-document)
SentRE [161] 2019 RE (Sentence-level)
Curriculum-RE [162] 2021 RE (Sentence-level)
DREEAM [163] 2023 RE (document-level)
Kumar et al. [97] 2020 End-to-End Construction
Guo et al. [164] 2021 End-to-End Construction
Grapher [41] 2021 End-to-End Construction
PiVE [165] 2023 End-to-End Construction
COMET [166] 2019 Distilling KGs from LLMs
BertNet [167] 2022 Distilling KGs from LLMs
West et al. [168] 2022 Distilling KGs from LLMs
LLM-augmented KG-to-text Generation Ribeiro et al [169] 2021 Leveraging Knowledge from LLMs
JointGT [42] 2021 Leveraging Knowledge from LLMs
FSKG2Text [170] 2021 Leveraging Knowledge from LLMs
GAP [171] 2022 Leveraging Knowledge from LLMs
GenWiki [172] 2020 Constructing KG-text aligned Corpus
KGPT [173] 2020 Constructing KG-text aligned Corpus
LLM-augmented KGQA Luo et al. [174] 2020 Entity/Relation Extractor
QA-GNN [117] 2021 Entity/Relation Extractor
Lukovnikov et al. [175] 2023 Entity/Relation Extractor
Nan et al. [176] 2023 Entity/Relation Extractor
DEKCOR [177] 2021 Answer Reasoner
DRLK [178] 2022 Answer Reasoner
OreoLM [179] 2022 Answer Reasoner
GreaseLM [119] 2022 Answer Reasoner
ReLMKG [180] 2022 Answer Reasoner
UniKGQA [43] 2023 Answer Reasoner
5 LLM-augmented for KGs
Knowledge graphs are famous for representing knowledge in a structural manner. They have been applied in many downstream tasks such as question answering, recommendation, and web search. However, the conventional KGs are often incomplete and existing methods often lack considering textual information. To address these issues, recent research has explored integrating LLMs to augment KGs to consider the textual information and improve the performance in downstream tasks. In this section, we will introduce the recent research on LLM-augmented KGs. Representative works are summarized in Table III. We will introduce the methods that integrate LLMs for KG embedding, KG completion, KG construction, KG-to-text generation, and KG question answering, respectively.
@@ 5.1 LLM-augmented KG Embedding
Knowledge graph embedding (KGE) aims to map each entity and relation into a low-dimensional vector (embedding) space. These embeddings contain both semantic and structural information of KGs, which can be utilized for various tasks such as question answering [181], reasoning [38], and recommendation [182]. Conventional knowledge graph embedding methods mainly rely on the structural information of KGs to optimize a scoring function defined on embeddings (e.g., TransE [25], and DisMult [183]). However, these approaches often fall short in representing unseen entities and long-tailed relations due to their limited structural connectivity [184, 185]. To address this issue, as shown in Fig. 16, recent research adopt LLMs to enrich representations of KGs by encoding the textual descriptions of entities and relations [96, 40].
@@ 5.1.1 LLMs as Text Encoders
Figure 16: LLMs as text encoder for knowledge graph embedding (KGE).
@@ 5.2 LLM-augmented KG Completion
As shown in Fig. 18 (a), (b), and (c), This line of work first uses encoder-only LLMs to encode textual information as well as KG facts. Then, they predict the plausibility of the triples by feeding the encoded representation into a prediction head, which could be a simple MLP or conventional KG score function (e.g., TransE [25] and TransR [186]).
@@
5.2.3 Comparison between PaE and PaG
LLMs as Encoders (PaE) applies an additional prediction head on the top of the representation encoded by LLMs. Therefore, the PaE framework is much easier to finetune since we can only optimize the prediction heads and freeze the LLMs. Moreover, the output of the prediction can be easily specified and integrated with existing KGC functions for different KGC tasks. However, during the inference stage, the PaE requires to compute a score for every candidate in KGs, which could be computationally expensive. Besides, they cannot generalize to unseen entities. Furthermore, the PaE requires the representation output of the LLMs, whereas some state-of-the-art LLMs (e.g. GPT-411footnotemark: 1) are closed sources and do not grant access to the representation output.
@@ 5.3.3 Relation Extraction (RE)
5.3.4 End-to-End KG Construction
Currently, researchers are exploring the use of LLMs for end-to-end KG construction. Kumar et al. [97] propose a unified approach to build KGs from raw text, which contains two LLMs powered components. They first finetune a LLM on named entity recognition tasks to make it capable of recognizing entities in raw text. Then, they propose another “2-model BERT” for solving the relation extraction task, which contains two BERT-based classifiers. The first classifier learns the relation class whereas the second binary classifier learns the direction of the relations between the two entities. The predicted triples and relations are then used to construct the KG. Guo et al. [164] propose an end-to-end knowledge extraction model based on BERT, which can be applied to construct KGs from Classical Chinese text. Grapher [41] presents a novel end-to-end multi-stage system. It first utilizes LLMs to generate KG entities, followed by a simple relation construction head, enabling efficient KG construction from the textual description. PiVE [165] proposes a prompting with an iterative verification framework that utilizes a smaller LLM like T5 to correct the errors in KGs generated by a larger LLM (e.g., ChatGPT). To further explore advanced LLMs, AutoKG design several prompts for different KG construction tasks (e.g., entity typing, entity linking, and relation extraction). Then, it adopts the prompt to perform KG construction using ChatGPT and GPT-4.
@@ 5.5.2 LLMs as Answer Reasoners
Answer reasoners are designed to reason over the retrieved facts and generate answers. LLMs can be used as answer reasoners to generate answers directly. For example, as shown in Fig. 3 24, DEKCOR [177] concatenates the retrieved facts with questions and candidate answers as
@@
6 Synergized LLMs + KGs
The synergy of LLMs and KGs has attracted increasing attention these years, which marries the merits of LLMs and KGs to mutually enhance performance in various downstream applications. For example, LLMs can be used to understand natural language, while KGs are treated as a knowledge base, which provides factual knowledge. The unification of LLMs and KGs could result in a powerful model for knowledge representation and reasoning. In this section, we will discuss the Synergized LLMs + KGs from two perspectives: 1) knowledge representation, and 2) reasoning. We have summarized the representative works in Table IV.
@@ 6 Synergized LLMs + KGs
Knowledge representation JointGT [42] 2021
@@
RecInDial [243] 2022
KnowledgeDA [244] 2022
6.1 Knowledge Representation
Text corpus and knowledge graphs both contain enormous knowledge. However, the knowledge in the text corpus is usually implicit and unstructured, while the knowledge in KGs is explicit and structured. Therefore, it is necessary to align the knowledge in the text corpus and KGs to represent them in a unified way. The general framework of unifying LLMs and KGs for knowledge representation is shown in Fig. 25.
@@ 6.1 Knowledge Representation
6.2 Reasoning
To take advantage of both LLMs and KGs, researchers synergize LLMs and KGs to perform reasoning on various applications. In the question answering task, QA-GNN [117] first utilizes LLMs to process the text question and guide the reasoning step on the KGs. In this way, it can bridge the gap between text and structural information, which provides interpretability for the reasoning process. In the knowledge graph reasoning task, LARK [45] proposes a LLM-guided logical reasoning method. It first transforms the conventional logical rules into a language sequence and then asks LLMs to reason the final outputs. Moreover, Siyuan et al. [46] unify structure reasoning and language mode pre-training in a unified framework. Given a text input, they adopt LLMs to generate the logical query, which is executed on the KGs to obtain structural context. Last, the structural context is fused with textual information to generate the final output. RecInDial [243] combines the knowledge graphs and LLMs to provide personalized recommendations in the dialogue system. KnowledgeDA [244] proposes a unified domain language model development pipline to enhance the task-specific training procedure with domain knowledge graphs.
Figure 25: The general framework of unifying LLMs and KGs for knowledge representation.
7 Future Directions
In the previous sections, we have reviewed the recent advances in unifying KGs and LLMs, but here are still many challenges and open problems that need to be addressed. In this section, we discuss the future directions of this research area.
@@ 7.2 KGs for Editing Knowledge in LLMs
Although LLMs are capable of storing massive real-world knowledge, they cannot quickly update their internal knowledge updated as real-world situations change. There are some research efforts proposed for editing knowledge in LLMs [248, 249] without re-training the whole LLMs. Yet, such solutions still suffer from poor performance or computational overhead. Existing studies [250, 251] also propose solutions to edit knowledge in LLMs, but only limit themselves to handling simple tuple-based knowledge in KGs. In addition, there are still challenges, such as catastrophic forgetting and incorrect knowledge editing [252], leaving much room for further research.

### v3

@@ 4.4.1 KGs for LLM Probing
However, LAMA ignores the fact that the prompts are inappropriate. For example, the prompt “Obama worked as a _” may be more favorable to the prediction of the blank by the language models than “Obama is a _ by profession”. Thus, LPAQA [118] proposes a mining and paraphrasing-based method to automatically generate high-quality and diverse prompts for a more accurate assessment of the knowledge contained in the language model. Moreover, Adolphs et al. [128] attempt to use examples to make the language model understand the query, and experiments obtain substantial improvements for BERT-large on the T-REx data. Unlike using manually defined prompt templates, Autoprompt [119] proposes an automated method, which is based on the gradient-guided search to create prompts. LLM-facteval [121] designs a systematic framework that automatically generates probing questions from KGs. The generated questions are then used to evaluate the factual knowledge stored in LLMs.
@@ 4.4.2 KGs for LLM Analysis
5 LLM-augmented KGs
Knowledge graphs are famous for representing knowledge in a structural manner. They have been applied in many downstream tasks such as question answering, recommendation, and web search. However, the conventional KGs are often incomplete and existing methods often lack considering textual information. To address these issues, recent research has explored integrating LLMs to augment KGs to consider the textual information and improve the performance in downstream tasks. In this section, we will introduce the recent research on LLM-augmented KGs. We will introduce the methods that integrate LLMs for KG embedding, KG completion, KG construction, KG-to-text generation, and KG question answering, respectively. Representative works are summarized in Table III.
@@ 5 LLM-augmented KGs
[b] Task Method Year LLM Technique LLM-augmented KG embedding Pretrain-KGE [94] 2020 E LLMs as Text Encoders KEPLER [40] 2020 E LLMs as Text Encoders Nayyeri et al. [132] 2022 E LLMs as Text Encoders Huang et al. [133] 2022 E LLMs as Text Encoders CoDEx [134] 2022 E LLMs as Text Encoders LMKE [135] 2022 E LLMs for Joint Text and KG Embedding kNN-KGE [136] 2022 E LLMs for Joint Text and KG Embedding LambdaKG [137] 2023 E + D + ED LLMs for Joint Text and KG Embedding LLM-augmented KG completion KG-BERT [26] 2019 E Joint Encoding MTL-KGC [138] 2020 E Joint Encoding PKGC [139] 2022 E Joint Encoding LASS [140] 2022 E Joint Encoding MEM-KGC [141] 2021 E MLM Encoding OpenWorld KGC [142] 2023 E MLM Encoding StAR [143] 2021 E Separated Encoding SimKGC [144] 2022 E Separated Encoding LP-BERT [145] 2022 E Separated Encoding GenKGC [96] 2022 ED LLM as decoders KGT5 [146] 2022 ED LLM as decoders KG-S2S [147] 2022 ED LLM as decoders AutoKG [93] 2023 D LLM as decoders LLM-augmented KG construction ELMO [148] 2018 E Named Entity Recognition GenerativeNER [149] 2021 ED Named Entity Recognition LDET [150] 2019 E Entity Typing BOX4Types [151] 2021 E Entity Typing ELQ [152] 2020 E Entity Linking ReFinED [153] 2022 E Entity Linking BertCR [154] 2019 E CR (Within-document) Spanbert [155] 2020 E CR (Within-document) CDLM [156] 2021 E CR (Cross-document) CrossCR [157] 2021 E CR (Cross-document) CR-RL [158] 2021 E CR (Cross-document) SentRE [159] 2019 E RE (Sentence-level) Curriculum-RE [160] 2021 E RE (Sentence-level) DREEAM [161] 2023 E RE (Document-level) Kumar et al. [95] 2020 E End-to-End Construction Guo et al. [162] 2021 E End-to-End Construction Grapher [41] 2021 ED End-to-End Construction PiVE [163] 2023 D + ED End-to-End Construction COMET [164] 2019 D Distilling KGs from LLMs BertNet [165] 2022 E Distilling KGs from LLMs West et al. [166] 2022 D Distilling KGs from LLMs LLM-augmented KG-to-text Generation Ribeiro et al [167] 2021 ED Leveraging Knowledge from LLMs JointGT [42] 2021 ED Leveraging Knowledge from LLMs FSKG2Text [168] 2021 D + ED Leveraging Knowledge from LLMs GAP [169] 2022 ED Leveraging Knowledge from LLMs GenWiki [170] 2020 - Constructing KG-text aligned Corpus KGPT [171] 2020 ED Constructing KG-text aligned Corpus LLM-augmented KGQA Lukovnikov et al. [172] 2019 E Entity/Relation Extractor Luo et al. [173] 2020 E Entity/Relation Extractor QA-GNN [131] 2021 E Entity/Relation Extractor Nan et al. [174] 2023 E + D + ED Entity/Relation Extractor DEKCOR [175] 2021 E Answer Reasoner DRLK [176] 2022 E Answer Reasoner OreoLM [177] 2022 E Answer Reasoner GreaseLM [178] 2022 E Answer Reasoner ReLMKG [179] 2022 E Answer Reasoner UniKGQA [43] 2023 E Answer Reasoner
  • E: Encoder-only LLMs, D: Decoder-only LLMs, ED: Encoder-decoder LLMs.
@@ 5.1 LLM-augmented KG Embedding
Knowledge graph embedding (KGE) aims to map each entity and relation into a low-dimensional vector (embedding) space. These embeddings contain both semantic and structural information of KGs, which can be utilized for various tasks such as question answering [180], reasoning [38], and recommendation [181]. Conventional knowledge graph embedding methods mainly rely on the structural information of KGs to optimize a scoring function defined on embeddings (e.g., TransE [33], and DisMult [182]). However, these approaches often fall short in representing unseen entities and long-tailed relations due to their limited structural connectivity [183, 184]. To address this issue, as shown in Fig. 14, recent research adopts LLMs to enrich representations of KGs by encoding the textual descriptions of entities and relations [94, 40].
Figure 14: LLMs as text encoder for knowledge graph embedding (KGE).
@@ 5.2 LLM-augmented KG Completion
As shown in Fig. 16 (a), (b), and (c), this line of work first uses encoder-only LLMs to encode textual information as well as KG facts. Then, they predict the plausibility of the triples or masked entities by feeding the encoded representation into a prediction head, which could be a simple MLP or conventional KG score function (e.g., TransE [33] and TransR [185]).
@@
Figure 18: The framework of prompt-based PaG for KG Completion.
Comparison between PaE and PaG. LLMs as Encoders (PaE) applies an additional prediction head on the top of the representation encoded by LLMs. Therefore, the PaE framework is much easier to finetune since we can only optimize the prediction heads and freeze the LLMs. Moreover, the output of the prediction can be easily specified and integrated with existing KGC functions for different KGC tasks. However, during the inference stage, the PaE requires to compute a score for every candidate in KGs, which could be computationally expensive. Besides, they cannot generalize to unseen entities. Furthermore, the PaE requires the representation output of the LLMs, whereas some state-of-the-art LLMs (e.g. GPT-411footnotemark: 1) are closed sources and do not grant access to the representation output.
@@ 5.3.3 Relation Extraction (RE)
End-to-End KG Construction. Currently, researchers are exploring the use of LLMs for end-to-end KG construction. Kumar et al. [95] propose a unified approach to build KGs from raw text, which contains two LLMs powered components. They first finetune a LLM on named entity recognition tasks to make it capable of recognizing entities in raw text. Then, they propose another “2-model BERT” for solving the relation extraction task, which contains two BERT-based classifiers. The first classifier learns the relation class whereas the second binary classifier learns the direction of the relations between the two entities. The predicted triples and relations are then used to construct the KG. Guo et al. [162] propose an end-to-end knowledge extraction model based on BERT, which can be applied to construct KGs from Classical Chinese text. Grapher [41] presents a novel end-to-end multi-stage system. It first utilizes LLMs to generate KG entities, followed by a simple relation construction head, enabling efficient KG construction from the textual description. PiVE [163] proposes a prompting with an iterative verification framework that utilizes a smaller LLM like T5 to correct the errors in KGs generated by a larger LLM (e.g., ChatGPT). To further explore advanced LLMs, AutoKG design several prompts for different KG construction tasks (e.g., entity typing, entity linking, and relation extraction). Then, it adopts the prompt to perform KG construction using ChatGPT and GPT-4.
@@ 5.5.2 LLMs as Answer Reasoners
Answer reasoners are designed to reason over the retrieved facts and generate answers. LLMs can be used as answer reasoners to generate answers directly. For example, as shown in Fig. 3 22, DEKCOR [175] concatenates the retrieved facts with questions and candidate answers as
@@
Synergized Knowledge representation JointGT [42] 2021
@@
KSL [239] 2023
StructGPT [237] 2023
Think-on-graph [240] 2023
6 Synergized LLMs + KGs
The synergy of LLMs and KGs has attracted increasing attention these years, which marries the merits of LLMs and KGs to mutually enhance performance in various downstream applications. For example, LLMs can be used to understand natural language, while KGs are treated as a knowledge base, which provides factual knowledge. The unification of LLMs and KGs could result in a powerful model for knowledge representation and reasoning.
In this section, we will discuss the state-of-the-art Synergized LLMs + KGs from two perspectives: 1) Synergized Knowledge Representation, and 2) Synergized Reasoning. Representative works are summarized in Table IV.
Figure 23: Synergized knowledge representation by additional KG fusion modules.
6.1 Synergized Knowledge Representation
Text corpus and knowledge graphs both contain enormous knowledge. However, the knowledge in text corpus is usually implicit and unstructured, while the knowledge in KGs is explicit and structured. Synergized Knowledge Representation aims to design a synergized model that can effectively represent knowledge from both LLMs and KGs. The synergized model can provide a better understanding of the knowledge from both sources, making it valuable for many downstream tasks.
To jointly represent the knowledge, researchers propose the synergized models by introducing additional KG fusion modules, which are jointly trained with LLMs. As shown in Fig. 23, ERNIE [35] proposes a textual-knowledge dual encoder architecture where a T-encoder first encodes the input sentences, then a K-encoder processes knowledge graphs which are fused them with the textual representation from the T-encoder. BERT-MK [241] employs a similar dual-encoder architecture but it introduces additional information of neighboring entities in the knowledge encoder component during the pre-training of LLMs. However, some of the neighboring entities in KGs may not be relevant to the input text, resulting in extra redundancy and noise. CokeBERT [242] focuses on this issue and proposes a GNN-based module to filter out irrelevant KG entities using the input text. JAKET [243] proposes to fuse the entity information in the middle of the large language model.
@@ 6.1 Synergized Knowledge Representation
6.2 Synergized Reasoning
To better utilize the knowledge from text corpus and knowledge graph reasoning, Synergized Reasoning aims to design a synergized model that can effectively conduct reasoning with both LLMs and KGs.
LLM-KG Fusion Reasoning. LLM-KG Fusion Reasoning leverages two separated LLM and KG encoders to process the text and relevant KG inputs [244]. These two encoders are equally important and jointly fusing the knowledge from two sources for reasoning. To improve the interaction between text and knowledge, KagNet [38] proposes to first encode the input KG, and then augment the input textual representation. In contrast, MHGRN [234] uses the final LLM outputs of the input text to guide the reasoning process on the KGs. Yet, both of them only design a single-direction interaction between the text and KGs. To tackle this issue, QA-GNN [131] proposes to use a GNN-based model to jointly reason over input context and KG information via message passing. Specifically, QA-GNN represents the input textual information as a special node via a pooling operation and connects this node with other entities in KG. However, the textual inputs are only pooled into a single dense vector, limiting the information fusion performance. JointLK [245] then proposes a framework with fine-grained interaction between any tokens in the textual inputs and any KG entities through LM-to-KG and KG-to-LM bi-directional attention mechanism. As shown in Fig. 24, pairwise dot-product scores are calculated over all textual tokens and KG entities, the bi-directional attentive scores are computed separately. In addition, at each jointLK layer, the KGs are also dynamically pruned based on the attention score to allow later layers to focus on more important sub-KG structures. Despite being effective, in JointLK, the fusion process between the input text and KG still uses the final LLM outputs as the input text representations. GreaseLM [178] designs deep and rich interaction between the input text tokens and KG entities at each layer of the LLMs. The architecture and fusion approach is mostly similar to ERNIE [35] discussed in Section 6.1, except that GreaseLM does not use the text-only T-encoder to handle input text.
Figure 24: The framework of LLM-KG Fusion Reasoning.
LLMs as Agents Reasoning. Instead using two encoders to fuse the knowledge, LLMs can also be treated as agents to interact with the KGs to conduct reasoning [246], as illustrated in Fig. 25. KD-CoT [247] iteratively retrieves facts from KGs and produces faithful reasoning traces, which guide LLMs to generate answers. KSL [239] teaches LLMs to search on KGs to retrieve relevant facts and then generate answers. StructGPT [237] designs several API interfaces to allow LLMs to access the structural data and perform reasoning by traversing on KGs. Think-on-graph [240] provides a flexible plug-and-play framework where LLM agents iteratively execute beam searches on KGs to discover the reasoning paths and generate answers. To enhance the agent abilities, AgentTuning [248] presents several instruction-tuning datasets to guide LLM agents to perform reasoning on KGs.
Figure 25: Using LLMs as agents for reasoning on KGs.
Comparison and Discussion. LLM-KG Fusion Reasoning combines the LLM encoder and KG encoder to represent knowledge in a unified manner. It then employs a synergized reasoning module to jointly reason the results. This framework allows for different encoders and reasoning modules, which are trained end-to-end to effectively utilize the knowledge and reasoning capabilities of LLMs and KGs. However, these additional modules may introduce extra parameters and computational costs while lacking interpretability. LLMs as Agents for KG reasoning provides a flexible framework for reasoning on KGs without additional training cost, which can be generalized to different LLMs and KGs. Meanwhile, the reasoning process is interpretable, which can be used to explain the results. Nevertheless, defining the actions and policies for LLM agents is also challenging. The synergy of LLMs and KGs is still an ongoing research topic, with the potential to have more powerful frameworks in the future.
7 Future Directions and Milestones
In this section, we discuss the future directions and several milestones in the research area of unifying KGs and LLMs.
@@ 7.2 KGs for Editing Knowledge in LLMs
Although LLMs are capable of storing massive real-world knowledge, they cannot quickly update their internal knowledge updated as real-world situations change. There are some research efforts proposed for editing knowledge in LLMs [252] without re-training the whole LLMs. Yet, such solutions still suffer from poor performance or computational overhead [253]. Existing studies [254] also reveal that edit a single fact would cause a ripple effect for other related knowledge. Therefore, it is necessary to develop a more efficient and effective method to edit knowledge in LLMs. Recently, researchers try to leverage KGs to edit knowledge in LLMs efficiently.